OpenStudy (anonymous):

Given the system of equations, what is the solution? 3x - 2y + 10 = 0 5y = 4x + 8 {(-34/7, 16/7)} {(34/7, -16/7)} {(-34/7, -16/7)}

OpenStudy (anonymous):

Substitute and check.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

a solution is an order pair (x, y) that satisfy both equations in the system... so if you want to determine if an ordered pair is a solution, take one of those choices and plug them into both equations...

OpenStudy (anonymous):

\[\left(-\frac{34}{7},-\frac{16}{7}\right) \]